# MOSAIC: Modeling Social AI for Content Dissemination and Regulation in Multi-Agent Simulations

### v1

@@ H.1 Power-Law Distribution of User Popularity
First we define the popularity of users as a sum of the number of followers, number of likes, shares, and comments received. We collected the top 50 users and plotted their popularity (as measured by the sum of engagement received by them) from highest to lowest in Figure 5. We observe a power-law distribution of user influence. We have $f(x)=120x^{-0.6}$ as the best-fitting power-law approximation of our sampled data, as shown in Figure 6. With $\alpha=0.60$, our regression line has an $R^{2}=0.84$, suggesting a strong fit to our user engagement data and that our social system follows a typical power law distribution where a few users generate most of the engagement. Existing analysis on real-world social networks suggest that this power-law exponent usually ranges from 1.5-2.5 depending on the specific context (Muchnik et al., 2013; Bild et al., 2015). Our best-fit exponent is lower than these reported numbers, but it still illustrates a clear trend that a minority of users/content collect most of the engagement while the majority of them do not contribute nearly as much.
@@
H.2 Persona Attributes Don’t Correlate with Engagement
We analyzed user engagement by comparing the top 50 most engaged users (highest number of followers, likes, shares, comments, etc.) with the bottom 50 least engaged users across several attributes. The Chi-square test results summarized in Table 3 indicate that there are no statistically significant differences in the distributions of age group, gender, activity type, hobby, ethnicity, income level, political affiliation, or primary goal between the two groups. Although some attributes, such as ethnicity and hobby, exhibited medium effect sizes (Cramer’s V Cramér (1946) of 0.319 and 0.302, respectively), their associated p-values did not reach conventional levels of statistical significance. This suggests that the attributes examined do not notably influence the level of user engagement.
@@ H.2 Persona Attributes Don’t Correlate with Engagement
H.3 Do Content Topics Matter?
@@
H.4 Clues from Agents’ Own Reasoning Traces and Recommender System
@@
The analysis of agent reasoning reveals patterns in how agents engage with content and users on social media. As shown in Figure 4, agents demonstrate clear and distinct emotional sentiment patterns associated with different types of actions. Positive-dominant actions such as following users (99% positive sentiment), commenting (97%), liking posts (92%), and sharing content (92%) indicate that agents predominantly perceive their interactions as constructive contributions. Conversely, negative sentiment predominantly characterizes actions like flagging posts (71% negative) and unfollowing users (40% negative), reflecting agents’ use of these interactions primarily for expressing disapproval or concern.
@@
Appendix I Prompts
I.1 Post Creation Prompt
@@
I.2 Feed Reaction Prompt
@@ 5. Goals or objectives you might want to pursue
I.4 Fact-checking Prompt

### v2

@@ A.2 Power-Law Distribution of User Popularity
First we define the popularity of users as a sum of the number of followers, number of likes, shares, and comments received. We collected the top 50 users and plotted their popularity (as measured by the sum of engagement received by them) from highest to lowest in Fig. 8. We observe a power-law distribution of user influence. We have $f(x)=120x^{-0.6}$ as the best-fitting power-law approximation of our sampled data, as shown in Fig. 5. With $\alpha=0.60$, our regression line has an $R^{2}=0.84$, suggesting a strong fit to our user engagement data and that our social system follows a typical power law distribution where a few users generate most of the engagement. Existing analysis on real-world social networks suggest that this power-law exponent usually ranges from 1.5-2.5 depending on the specific context (Muchnik et al., 2013; Bild et al., 2015). Our best-fit exponent is lower than these reported numbers, but it still illustrates a clear trend that a minority of users/content collect most of the engagement while the majority of them do not contribute nearly as much.
@@
A.3 Persona Attributes Don’t Correlate with Engagement
We analyzed user engagement by comparing the top 50 most engaged users (highest number of followers, likes, shares, comments, etc.) with the bottom 50 least engaged users across several attributes. The Chi-square test results summarized in Tab. 2 indicate that there are no statistically significant differences in the distributions of age group, gender, activity type, hobby, ethnicity, income level, political affiliation, or primary goal between the two groups. Although some attributes, such as ethnicity and hobby, exhibited medium effect sizes (Cramer’s V Cramér (1946) of 0.319 and 0.302, respectively), their associated p-values did not reach conventional levels of statistical significance. This suggests that the attributes examined do not notably influence the level of user engagement.
@@ A.3 Persona Attributes Don’t Correlate with Engagement
A.4 Do Content Topics Matter?
@@
A.5 Clues from Agents’ Own Reasoning Traces and Recommender System
@@
The analysis of agent reasoning reveals patterns in how agents engage with content and users on social media. As shown in Fig. 3, agents demonstrate clear and distinct emotional sentiment patterns associated with different types of actions. Positive-dominant actions such as following users (99% positive sentiment), commenting (97%), liking posts (92%), and sharing content (92%) indicate that agents predominantly perceive their interactions as constructive contributions. Conversely, negative sentiment predominantly characterizes actions like flagging posts (71% negative) and unfollowing users (40% negative), reflecting agents’ use of these interactions primarily for expressing disapproval or concern.
@@
Appendix B Persona Generation Details
B.1 Persona Replication from Human Survey
B.2 Synthetic Persona from Agent Bank
B.3 User Generation and Instantiation
Persona Generation
As illustrated in Fig. 1, personas are generated using a combination of predefined questions and sampling from probabilistic distributions stored in an agent_bank profile collection, inspired by Park et al. (2024). Key demographic attributes such as age, gender, ethnicity, and primary language are assigned probabilistically to mirror real-world distributions. For instance, age follows a normal distribution centered around 35 years, while other attributes are sampled based on predefined probabilities. We disclose these questions and describe more detail of the methodology in Appendix B. After synthesizing the structured profiles for agents, we construct a natural language description for each of them. This process leverages a mixture of deterministic rules and LLM-based augmentation using GPT-4o Hurst et al. (2024) to enhance diversity and realism.
User Instantiation
Once personas are generated, they are instantiated as agent users within the simulation. Each agent is assigned a unique user ID and a persona profile that includes background details and interest labels. The relational database serves as the backbone for recording agent activities, ensuring persistent storage of interactions, post engagements, and behavioral updates. This database facilitates dynamic user tracking and enables post-simulation analysis of engagement trends and content spread. We provide more implementation details in Appendix F.
Appendix C Models And Computational Budget
We use APIs for proprietary LLM inferences in our experiments. We locally hosted some smaller open-weight LLMs in the early exploratory stages of the projects, on H100 GPUs. We end up consuming approximately \$600 of OpenAI credits, \$300 of Anthropic API, and \$40 of DeepSeek API.
Appendix D Details of the Human Study and IRB
The study was open to 20,240 eligible participants from a larger Prolific population of 232,330, and we collected 204 valid responses from eligible participants. The survey was conducted via Prolific to collect responses from U.S.-based participants fluent in English. Participants were asked to complete a 12-minute survey assessing their demographic characteristics and social media interactions. The survey, hosted on Google Forms, required no software downloads or special device features and was accessible via mobile, tablet, or desktop. Participant recruitment applied custom screening for language, political spectrum, vaccine opinion, and prior participation, ensuring a targeted sample. Responses were collected using Prolific ID via a question at the start of the form, and participants received a completion code upon finishing. Compensation was set at \$2.40 per participant, equivalent to \$12.00/hour, and submissions were manually reviewed before approval. The median completion time was approximately 14.5 minutes. All members on our research team have obtained IRB approval before the human study was conducted. Our study costs a total of \$480 for participant payment and \$160 of platform fee. Fig. 6 shows the complete breakdown of the 9 key demographic distributions of the 204 human participants. This study was performed under approval from the appropriate institutional ethics review board. Full IRB documentation will be made available upon request.
Figure 6: Demographic Distributions of Study Participants.
Category Significant Differences Non-Significant Differences
Age 25-34 (shares) 18-24, 35-44, 45-54, 55-64, 65-74
Gender Male (likes, shares) Female
Asian (comments)
D.1 Per-Demographic Attribute Engagement Pattern
We also analyzed the reaction patterns between human participants and persona-driven agents grouped by specific demographic attributes such as age, gender, income, ethnicity, etc, as shown in Tab. 4.
Overall, out of 52 demographic groups analyzed, 14 show significant differences in one or more engagement types, while 38 do not. The criteria for significance were based on a p-value of less than 0.05 in statistical comparisons. The results suggest that agents may be more adept at simulating the engagement patterns of "common" or more broadly represented demographic groups in LLM pretraining data, as indicated by the lack of significant differences in many of these groups. We find that out of 52 examined demographic subgroups, only 14 showed statistically significant differences (p < 0.05) in at least one engagement metric (Tab. 4). Notable discrepancies appeared in the 25–34 age group (shares) and several religious, ethnic, educational, income, and political categories. However, most demographic groups exhibited no significant differences, suggesting that agents simulate typical engagement behavior more accurately for demographics more prevalent in LLM training data.
Appendix E An Extended Version of Related Work
Behavioral Economics and Persuasian Games.
AI-Driven Social Simulations and Generative Agents.
Misinformation Spread and Fact-Checking Mechanisms.
Simulations as Tools for Policy and Platform Governance
Appendix F Database Schema
Appendix G Agent Action Space
Figure 7: Agent’s action space.
G.1 Agent Decision-Making Process
Feed Presentation
Memory and Reflection Module
We implement an AgentMemory module which manages the memory and reflection capabilities of each agent. Memories are categorized into interactions (e.g., past engagements) and reflections (high-level insights derived from past behaviors). Each piece of memory is assigned an importance score, which decays over time unless reinforced by further interactions. The decay function ensures that long-term behaviors emerge naturally based on experience. Please refer to Appendix I for more details of the Memory module.
Agent Decision-Making and Action Execution
Appendix H Detailed Experiment Configuration
Appendix I Details of the Memory Module
Memory relevance is computed as:
[…]
The decay factor is defined as:
$\text{Decay}=\max(0,\text{PrevDecay}-\alpha\Delta t)$
Importance Scoring.
  • Emotional: love, hate, angry, happy, sad
  • Action: achieved, failed, learned, discovered
  • Relationship: friend, follow, connect, share
  • Goal: objective, target, aim, purpose
Let $k$ be the number of keyword matches in the memory content. Then:
$\text{Importance}=\min(1.0,0.5+0.1k)$
Figure 8: Top 50 users with highest engagement.
Appendix J Total Engagement Comparison
Figure 9: Total Engagement Across All Models and Fact Check Modes
Figure 10: Average Number of Likes
Fig. 9 tracks average interactions with factual and misinformative posts across time steps 0-40 for three LLMs: Claude-3.7-Sonnet, DeepSeek-V3, and GPT-4o over four fact-checking types as defined in the experiment configurations: None, Community Based, Third Party, and Hybrid. Each subplot shows how user engagement with true and false content evolves.
J.1 Total Interactions: Claude-3.7-Sonnet
"None", "Community Based": There is little to no interaction with misinformative content. Interaction with factual posts decreases over time.
"Third Party": Initially, there is higher interaction with misinformative content than factual content. This reverses as interaction with misinformation decreases and with factual content increases.
"Hybrid": There is little to no interaction with misinformative content. There is steady increase in interaction with factual content.
J.2 Total Interactions: DeepSeek-V3
"None", "Community Based", "Hybrid": There is little interaction with misinfomrative content. Interaction with factual content starts off strongly and increases over time.
"Third Party": Initially, there is much higher interaction with misinformative content than factual content. Over time, interaction with misinformation decreases and with factual content increases, but interaction with misinformation is still close to factual content interaction.
J.3 Total Interactions: GPT-4o
"None", "Community Based": There is little interaction with both factual and misinformative content.
"Third Party", "Hybrid": There is some interaction with misinformative content, though less than the consistent and increasing interaction with factual content.
Figure 11: Average Number of Note Ratings
J.4 Cross Model Comparisons
"None": All models have very little interaction with misinformative content. DeepSeek is the only model that increases interaction with factual content, while the others decrease over time.
"Community Based": All models continue to have very little interaction with misinformative content. Again, DeepSeek is the only model that increases interaction with factual content over time.
"Third Party": All experiments intially start with a higher level of interaction with misinformation than interaction with factual information. This trend reverses across all models over time. However, the level of difference in interaction with misinformation and factual information is most pronounced in Claude. Other models have less differentiation.
"Hybrid": Claude has the lowest levels of interaction with misinformative content as well as the highest and most consistent levels of interaction with factual content. DeepSeek initially displays the same trend, while experiencing increasing misinformation engagment in later time steps. GPT-4o has constant interaction with different content types, though slightly higher with factual content.
J.5 Potential Claims
Claude-3.7-Sonnet: Under the "None" and "Community Based" settings, Claude shows almost no interaction with misinformation. However, interactions with factual content decline steadily over time. This suggests that in the absence of strong moderation signals, even factual content loses traction and user engagement drops.
With "Third Party" moderation, there is initially more engagement with misinformation than with factual content. Over time, this trend reverses: misinformation engagement declines while factual engagement increases. This shift implies that authoritative third-party intervention can realign attention toward accurate information.
Under the "Hybrid" setting, Claude exhibits the strongest performance. Misinformation is almost completely suppressed, while factual content steadily gains engagement throughout the simulation. This indicates that Claude, when acting as an agent, is highly responsive to layered, multi-source moderation and can maintain a sustained pro-factual trajectory when given comprehensive oversight.
DeepSeek-V3: DeepSeek behaves differently from Claude. In the "None," "Community Based," and "Hybrid" settings, it maintains low misinformation engagement while steadily increasing interactions with factual content. This pattern suggests DeepSeek may have a stronger default tendency toward promoting truthful content, even with minimal intervention.
Under the "Third Party" condition, however, the simulation starts with high misinformation engagement. Although this decreases over time and begins to converge with factual engagement, misinformation remains close in magnitude. Unlike Claude, DeepSeek does not show a strong corrective response to third-party fact-checking, suggesting lower sensitivity to external moderation.
GPT-4o: GPT shows low overall engagement with both factual and misinformative content under the "None" and "Community Based" regimes. This could reflect a more cautious content sharing dynamic when no clear verification cues are available.
Under both the "Third Party" and "Hybrid" settings, factual engagement consistently increases, while misinformation remains relatively low. Though the effect is less dramatic than with Claude, GPT demonstrates a stable alignment with accurate content in the presence of reliable fact-checking mechanisms, suggesting moderate responsiveness to external moderation.
Our findings may suggest that layered, hybrid fact-checking is most effective overall, especially for models that are more responsive to external moderation like Our findings may suggest that layered, hybrid fact-checking is most effective overall, especially for models that are more responsive to external moderation like Claude and GPT. In contrast, models like DeepSeek may require tailored or persistent strategies to achieve similar alignment outcomes. and GPT. In contrast, models like DeepSeek may require tailored or persistent strategies to achieve similar alignment outcomes.
Figure 12: Average Number of Notes
Appendix K Disaggregated Engagement Comparison
We’ve decided to further explore interactions of all models across misinformative and factual content through five allowed actions: comments, likes, note ratings, community notes, and shares.
K.1 Average Comments
Fig. 14 shows Average Comments across all models. In "Hybrid", comments on factual content steadily increase over time, while misinformation receives minimal and flat engagement. In "Community Based" and "None", misinformation remains low and nearly flat, while factual content is slightly higher. In "Third-Party", there is a steady increase in comments on factual content, while comments on misinformative content start high and decrease. Strong moderation (like hybrid) seems to promote public discussion of factual content while limiting commentary on misinformation.
Figure 13: Average Number of Shares
Figure 14: Average Number of Comments
K.2 Average Likes
Fig. 10 shows Average Likes across all models. Likes across all fact checking modes have high engagement over time for factual information, with minimal engagement with misinformation. For "Hybrid" and "Third Party" approaches, likes trend up, while for "Community-Based" and "No Fact Checking", likes trend down over time. Interestingly, while there is almost no engagement in other modes, there remains some likes for misinformation in "Third-Party". Hybrid moderation seems to foster the strongest positive sentiment toward factual content with lowest disengagement with misinformative content.
K.3 Average Note Ratings
Fig. 11 shows Average Note Rating count across all models. There are a higher amount of note ratings in "Hybrid" than "Community-Based". Additionally, expected behavior is that there are few notes to rate for factual information.
K.4 Average Community Notes
Fig. 12 shows Average Community Note count across all models. In "Hybrid", there is a steady growth of community notes for misinformation the majority of the time steps. In "Community-Based", there is actually a steady decline. There are also no notes for factual information in "Community-Based", while there are few but steadily growing notes in "Hybrid".
Note Ratings and Community Notes suggests that a "Hybrid" approach best stimulates robust feedback loops for users, including interaction with all types of information and interaction with other users’ notes.
K.5 Average Shares
Fig. 13 shows the final action possible for agents, Average Share count across all models. The number of average shares is generally low, being less than 1 per post. In this proportion, for "Hybrid", "Community-Based", and "None" approaches there is a fast increase to decrease and constant low in sharing of misinformation. "Hybrid" has gradually increasing shares for factual information, while "Community-Based" and "None" slowly decrease, though "None" has more overall shares. "Third-Party" experiences the most overall shares, and the sharpest increase in shares of factual information over time as well. Overall, strong fact-checking encourages users to amplify factual content, whereas misinformation rarely achieves viral spread across any setting.
Appendix L Prompts
L.1 Post Creation Prompt
@@ Appendix I Details of the Memory Module
L.2 Feed Reaction Prompt
@@ 5. Goals or objectives you might want to pursue
L.4 Fact-checking Prompt